OpenStudy (anonymous):

Subtract and write in simplest form:

OpenStudy (anonymous):

subtract and write what in simplest form?

OpenStudy (anonymous):

\[\frac{ x+8 }{ 7 } -\frac{ x+3 }{ 7 }\]

OpenStudy (anonymous):

Well you already have a common denominator, so just subtract.

OpenStudy (anonymous):

\[\huge \frac{ x+8-x+3 }{ 7 }\]

OpenStudy (anonymous):

When we have same denominators then the answer is the difference between the numerators (a – c) over their common denominator. that is x + 8 - (x + 3) x + 8 - x - 3 5 ------------------ = ------------ = ----- 7 7 7